# Nonclassical microwave radiation from the parametric dynamical Casimir effect in the reversed-dissipation regime of circuit optomechanics

### v1

@@ Abstract
We propose an experimentally feasible optomechanical system (OMS) that is dispersively driven and operates in the reversed dissipation regime (RDR), where the mechanical damping rate far exceeds the cavity decay rate. We demonstrate that coherent, fast-time modulation of the driving laser frequency—on time scales longer than the mechanical decoherence time—allows for adiabatic elimination of the mechanical mode, resulting in strong parametric amplification of quantum vacuum fluctuations of the intracavity field. This mechanism, known as the parametric dynamical Casimir effect (parametric-DCE), leads to the generation of Casimir photons. In the dispersive RDR, we find that the total system Hamiltonian—including the DCE term—is intrinsically modified by a generalized optomechanical Kerr-type nonlinearity. This nonlinearity not only saturates the mean number of radiated Casimir photons on short time scales, even without dissipation, but also induces oscillatory behavior in their dynamics and quantum characteristics. Remarkably, the presence of the Kerr nonlinearity causes the generated DCE photons to exhibit nonclassical features, including sub-Poissonian statistics, negative Wigner function and quadrature squeezing which can be controlled by adjusting the system parameters. The proposed nonclassical microwave radiation source possesses the potential to be applied in quantum information processing, quantum computing as well as microwave quantum sensing.
@@ I Introduction
It is well-known that quantum parametric processes provide essential mechanisms for generating and manipulating quantum states in diverse aspects of quantum science and technologies. The most remarkable property of these processes concerns the dynamical amplification of quantum vacuum fluctuations that is responsible for the creation of pairs of real particles as a consequence of strong non-adiabatic change of a system parameter or boundary condition Moore1970 ; Davies1976 . This effect, which is generally referred to as the dynamical Casimir effect (DCE) Moore1970 ; Yablonovitch1989 ; Schwinger1992 ; Nation2012 ; Dodonov2020 , describes a process in which a cavity with periodically oscillating mirrors produces pairs of photons out of the electromagnetic vacuum fluctuations. Along with the theoretical studies on the issue of the possibility of particle creation via the DCE in a large variety of systems, ranging from cosmology Brevik2000 ; Durrer2007 ; Lock2017 to non-stationary cavity QED Dalvit1999 ; Crocce2002 ; Dodonov2003 ; Dodonov2011 ; Mendoza2019 ; Lo2020 , various theoretical schemes for practical applications of the DCE have been suggested, including generation of photons with nonclassical properties Dodonov1990 ; Dodonov1999 ; Johansson2013 , generation of squeezed Aggarwal2015 , and entangled atomic states Lange2018 , generation of multipartite entanglement in networks of quantum cavities and superconducting devices Felicetti2014 , generation of entanglement between two moving qubits Agusti2019 , generation of EPR quantum steering and Gaussian interferometric power Sabin2015 , and quantum synchronization of two qubits Mitarai2024 . In addition, in the context of quantum thermodynamics, it has been shown that the DCE can be used for realizing the so-called anti-DCE Sousa2015 , i.e., coherent annihilation of photons instead of creation, as a resource for work extraction from the atom-field system Dodonov2017 , for implementation of a quantum Otto heat engine Ferreri2023 , and for cooling down the cavity wall in the presence of a non-vanishing temperature gradient between the wall and the cavity Ferreri2024 .
@@
From an experimental point of view, for a measurable flux of real photons to be generated, the moving boundaries should oscillate at very high frequencies (more than 10GHz) that are not yet experimentally achievable. Consequently, alternative schemes based on imitation of boundary motion have been proposed, with examples including periodic modulation of the optical properties of the boundary Lombardi ; Dodonov2006 ; Dodonov2006-2 or of the optical path length of a cavity Dezael ; Faccio ; Motazedifard DCE , amplification of the DCE in a cavity within the Rabi model in the ultrastrong coupling regime via quantum optimal control strategies Hoeb , and time modulation of the Kerr or higher-order nonlinearities in a cavity Dodonov2022 . Some other experimental schemes aiming at the observation of the DCE can be found in Refs. Lombardi ; Agnesi ; Kawakubo . Notably, it has been reported about a successful implementation of DCE in superconducting circuits through two independent experiments by fast-modulating either the electrical boundary condition of a transmission line Pourkabirian or the effective speed of light in a Josephson metamaterial Lahteenmaki . Moreover, recently an analogue DCE has been experimentally realized in the near infrared regime using a dispersion-oscillating photonic crystal fiber vezzoli2019 . Results of this experiment implied that the generated DCE photons exhibit nonclassical photon anti-bunching statistics.
@@
Besides the photonic DCE, analog models for the generation of other particles than photons via the DCE have also been investigated such as the DCE of phonons in a time-modulated atomic Bose-Einstein condensate (BEC) Recati ; Jaskula , emission of Bogoliubov excitations in an exciton-polariton condensate by an ultrashort laser pulse Koghee , phononic DCE in a time-modulated quantum fluid of light Busch , DCE of magnon excitation in a spinor BEC driven by a time-dependent magnetic field Saito , and DCE of phonons in a gas of laser-cooled atoms with time-dependent effective charge Dodonov2014 .
@@
On the other hand, during the recent decades optomechanical systems (OMSs) where electromagnetic radiation pressure is linearly or quadratically coupled to a mechanical oscillator (MO) Aspelmeyer2014 have been widely employed in a large variety of applications, for example, displacement and force sensing Kippenberg2007 ; Tsang2010 ; Tsang2012 ; Bariani2015 ; Wimmer2014 ; Motazedifard2016 ; Allahverdi2022 ; Motazedifard2021 ; Motazedifard2019 ; Ebrahimi2021 ; Bemani2022 , optomechanical cooling Connell2010 ; Teufel2011 ; Chan2011 , quantum correlations such as entanglement Palomaki2013 ; Paternostro2007 ; Genes2009 ; Dalafi2018 ; Bemani2019 and synchronization Mari2013 ; Ying2014 ; Bemani2017 , generation of nonclassical states of the mechanical and optical modes Nunnenkamp2011 ; Jahne2009 ; Latmiral2018 ; Groblacher2018 ; Brunelli2018 ; Solki2023 , quadrature squeezing/amplification Kippenberg2008 ; Massel2017 ; Ockeloen2018 ; Motazedifard2019-2 , optomechanically induced transparency Xiong2018 ; Mikaeili2022 ; Agarwal2010 ; Weis2010 ; Motazedifard2021-2 ; Safavi2011 ; Karuza2013 ; Kronwald2013 ; Motazedifard2023 ; Yang2019 , and quantum illumination radar Barzanjeh2015 ; Barzanjeh2020 ; Allahverdi2024 .
@@
In recent years, a variety of theoretical studies have identified the OMSs as promising platforms for realizing the DCE. To mention some of them, one could highlight the DCE of phonons in a non-stationary quantum-well assisted optomechanical cavity Mahajan2015 ; the DCE of phononic excitation in the so-called membrane-in-the-middle optomechanical cavity Motazedifard2017 ; the DCE of photons and mechanical-/Bogoliubov-type phonons in a modulated hybrid optomechanical cavity containing an atomic BEC Motazedifard2018 ; the DCE of photons in an optomechanical cavity under incoherent mechanical excitation Settineri2019 ; the DCE of photons in a squeezed-cavity-assisted OMS Qin2019 ; the DCE of photons in an optomechanical cavity interacting with a one-dimensional photonic crystal Tanaka2020 ; the DCE of phonons in the ultrastrong-coupling regime of optomechanics Minganti2024 ; and the DCE of photons in non-perturbative coupling regime of cavity optomechanics Macri2018 . Remarkably, in Ref. Macri2018 the authors have found that the standard resonance condition for the generation of DCE photons which requires the mechanical frequency to be, at least, twice the first cavity mode frequency does not need to be satisfied, provided that the coupling between the cavity field and the moving mirror is non-vanishing with respect to the optical and mechanical resonance frequencies. In addition, they have shown that the non-perturbative regime of DCE can give rise to the steady-state entanglement between the moving mirror and the cavity field.
@@
One of the most important characteristics of cavity OMSs is a type of intrinsic nonlinearity resulting from the radiation pressure coupling. This nonlinearity stems from the fact that in a typical cavity OMS the position of the MO modulates the resonance frequency of the cavity mode. In other words, the optical length of the cavity depends on the intensity of the intracavity field and, consequently, the optomechanical cavity can behave effectively as a nonlinear Kerr medium Gong2009 ; Aldana2013 . This inherent nonlinearity enables pondermotive squeezing of the cavity field Fabre1994 ; Mancini1994 , generation of mechanical cat states Bose1997 ; Bose1999 , intracavity cat-state generation Leghtas2015 photon blockade Rabl2011 ; Ling2023 , optical bistability Dorsel1983 ; Gozzini1985 , phonon–photon entanglement in the bistable regime Ghobadi2011 , and cooling an optomechanical resonator into a cat state of motion Hauer2023 .
@@
The above-mentioned cavity optomechanical phenomena, such as mechanical ground-sate cooling, entanglement, quadrature squeezing/amplification, and optomechanically induced transparency emerge in the so-called normal dissipation regime, in which the decay rate of the MO is typically significantly smaller compared to that of the cavity mode. However, recent investigations in the field of dissipation engineering have introduced a novel regime of dissipation in optomechanics, the so-called reversed-dissipation regime (RDR), where the roles of the cavity mode and the MO are interchanged Nunnenkamp2014 , i.e., the dissipation rate of the MO prominently exceeds the cavity linewidth. In this regime, the MO plays the advantageous role of an additional cold dissipative reservoir for the cavity mode Wang2013 ; Metelmann2014 . Recently significant development  in experimental implementations has made possible the realization of the RDR in various experimental setups, including superconducting circuits Sliwa2015 ; Lecocq2017 , microwave-cavity optomechanical systems Toth2017 , and mechanical resonators embedded by erbium ions Ohta2021 . In addition, a variety of applications have been proposed and investigated for this regime. Some examples include quantum-limited amplification and self-oscillation of photons Nunnenkamp2014 ; Toth2017 , entangled-photon generation Wang2013 , nonreciprocal optomechanically induced transparency and enhancing the ground-state cooling of a MO Zhang2024 , and broadband nonreciprocal and chiral photon transmission Chen2023 .
@@
Inspired by the above-mentioned investigations, in this paper, we propose a protocol to study the parametric DCE in an experimentally realized microwave cavity OMS Toth2017 which is dispersively driven and operates in the RDR. In such case, the degrees of freedom of the MO can be adiabatically eliminated from the system dynamics. We show that the coherent time modulation of the driving laser frequency leads to the parametric amplification of the quantum vacuum fluctuations of the intracavity field mode, resulting in the creation of microwave Casimir photons over time scales longer than the mechanical decoherence time. In addition, the adiabatic elimination of the mechanical mode will induce an effective nonlinear Kerr-type photon–photon interaction, which provides the feasibility of controllable manipulation of the quantum properties of the generated DCE photons. We analyze both analytically and numerically the effects of the induced Kerr nonlinearity as well as the modulation amplitude of the driving laser frequency on the dynamical behaviors of the mean number of the generated microwave Casimir photons, photon counting statistics, and photon quadrature squeezing. Within the short-time approximation, i.e., over time scales much shorter than the cavity decay time, we find a closed analytical expression for the system unitary time-evolution operator by using the Wei-Norman theorem Norman1964 , which is of the form of a generalized squeezing operator affected by the induced Kerr nonlinearity. The presence of the Kerr nonlinearity causes that the time evolution of the mean number of Casimir photons exhibits an oscillatory behavior whose amplitude decreases with increasing time. Regarding the nonclassical features of the generated DCE photons, we find that they can exhibit quadrature squeezing and sub-Poissonian statistics in the course of time evolution. In particular, the induced Kerr nonlinearity provides the possibility of oscillatory transfer of squeezing between the position and momentum quadratures of the generated Casimir photons at certain time intervals. Moreover, due to the Kerr nonlinearity, the DCE photons manifest sub-Poissonian statistics over time scales shorter than the natural oscillation period of the MO as well as nonclassical state.
@@
The paper is structured as follows. First, we introduce the physical system under consideration in Sec. II, and derive an effective system Hamiltonian in the RDR of cavity optomechanics when the cavity is dispersively driven by an external modulated field. The effects of the induced Kerr nonlinearity on the dynamics of the mean number, the quantum statistics, and the quadrature squeezing of the generated Casimir photons are, respectively, discussed in Secs. IIIIV and V. In Sec. VI we explore the signature of the nonclassical characteristic of the generated Casimir photons by calculating the corresponding Wigner function. In particular, we examine the impact of system dissipation on the temporal behavior of the Wigner negativity. A detailed discussion of the experimental requirements for implementation of the model is undertaken in Sec. VII. We summarize the conclusions of our work in Sec. VIII. Some detailed mathematical calculations can be found in the appendices.
@@ II Physical Model and The Effective System Hamiltonian
Figure 1: (Color online) (a) Schematic diagram of the considered microwave optomechanical circuit Toth2017 composed of two inductively coupled microwave LC resonators which are inductively coupled to a microwave feedline as a common bath. The two hybridized dark and bright modes of the circuit that are, respectively, used as the primary $\hat{a}$ and auxiliary $\hat{a}_{\rm aux}$ interact with the vibration of the suspended top electrode of a shared capacitor, acting as a MO with resonance frequency $\omega_{m}$. The auxiliary (primary) mode is driven by a classical laser field of frequency $\omega_{L}^{\rm aux}(\omega_{L})$ and amplitude $E_{L}^{\rm aux}(E_{L})$. (b) To prepare a cold, dissipative mechanical reservoir for microwave photons the dissipation rate $\gamma_{m}$ of the MO should be increased to match the much larger dissipation rate $\kappa$ of the primary mode having resonance frequency $\omega_{c}$. For this purpose, the auxiliary mode with resonance frequency $\omega_{\rm aux}$ and dissipation rate $\kappa_{\rm aux}$ is used to damp out the MO via optomechanical sideband cooling and, hence, prepare it as a strongly dissipative, cold reservoir for the primary mode. (c) A schematic overview of the process to derive Hamiltonian (26).
As schematically illustrated in Fig. 1(a), the system we have chosen to consider is a microwave-cavity optomechanical circuit, which has been previously presented in Ref. Toth2017 aiming at realizing a cold, dissipative mechanical reservoir for microwave photons. It is composed of two inductively coupled microwave LC resonators, both coupled inductively to a common microwave feedline, serving as a common bath, with external coupling strengths $\kappa_{1}^{\rm ex}$ and $\kappa_{2}^{\rm ex}$. One of the LC resonators contains a mechanically compliant capacitor that can be modeled as a single-mode MO. By making use of an electromagnetic mode, referred to as the auxiliary mode, the MO is damped out via optomechanical sideband cooling Schliesser2008 in order to prepare it as a strongly dissipative, cold reservoir for another electromagnetic mode which referred to as the primary mode.
In order to accomplish the RDR in this setup, the auxiliary mode needs to be cooled down with a rate much faster than the electromagnetic decay rate of the primary mode, requiring the microwave cavities to have extremely different decay rates. One successful approach to overcome this challenge is to engineer hybridized modes with intrinsically different decay rates arising from interference in the output channel [for details, see the Supplementary Information in Ref. Toth2017 ]. The resulting interaction Hamiltonian is given by Toth2017 ($\hbar=1$)
@@
where $\hat{b}$ is the annihilation operator for the mechanical mode with frequency $\omega_{m}$ and decay rate $\gamma_{m}$, $\hat{a}_{1}$ and $\hat{a}_{2}$ are the annihilation operators for the bare modes with frequency $\omega_{0}$, $J$ denotes the intermode coupling strength, and $g_{0}^{\prime}$ stands for the vacuum electromechanical coupling strength to the first bare mode. By introducing the symmetric and antisymmetric superpositions of the bare modes $\hat{a}_{s,a}=(\hat{a}_{1}\pm\hat{a}_{2})/\sqrt{2}$ Toth2017 the intermode coupling term can be diagonalized, and thus the interaction Hamiltonian is expressed as
@@
[…]
The hybridized modes $\hat{a}_{s,a}$ constitute a strongly coupled (bright) and a weakly coupled (dark) mode with the respective decay rates $\kappa_{s}^{\rm ex}=\kappa_{1}^{\rm ex}+\kappa_{2}^{\rm ex}$ and $|\kappa_{a}^{\rm ex}=\absolutevalue{\kappa_{1}^{\rm ex}-\kappa_{2}^{\rm ex}}$ |, which can be explained as the result of the interference of the bare-mode external coupling rates $\kappa_{1}^{\rm ex}$ and $\kappa_{2}^{\rm ex}$ with the output channel [see Ref. Toth2017 and its Supplementary Information for details]. In the case of nearly the same coupling rates for the bare modes ($\kappa_{1}^{\rm ex}\approx\kappa_{2}^{\rm ex}$) we have $\kappa_{s}^{\rm ex}\gg\kappa_{a}^{\rm ex}$ so that the antisymmetric (dark) mode is approximately decoupled from the common reservoir. Following Ref. Toth2017 , in subsequent discussions we refer the dark and the bright modes as the primary and the auxiliary modes, respectively, with respective resonance frequencies $\omega_{c}$ and $\omega_{\rm aux}$ and decay rates $\kappa$ and $\kappa_{\rm aux}$ [see Fig. 1(b)].

### v2

@@ Abstract
We propose an experimentally feasible optomechanical system (OMS) that is dispersively driven and operates in the reversed dissipation regime (RDR), where the mechanical damping rate far exceeds the cavity decay rate. We demonstrate that coherent, fast-time modulation of the driving laser frequency-on time scales longer than the mechanical decoherence time-allows for adiabatic elimination of the mechanical mode, resulting in strong parametric amplification of quantum vacuum fluctuations of the intracavity field. This mechanism, known as the parametric dynamical Casimir effect (parametric-DCE), leads to the generation of Casimir photons. In the dispersive RDR, we find that the total system Hamiltonian—including the DCE term—is intrinsically modified by a generalized optomechanical Kerr-type nonlinearity. This nonlinearity not only saturates the mean number of radiated Casimir photons on short time scales, even without dissipation, but also induces oscillatory behavior in their dynamics and quantum characteristics. Remarkably, the presence of the Kerr nonlinearity causes the generated DCE photons to exhibit nonclassical features, including simultaneous sub-Poissonian statistics and negative Wigner function, as well as quadrature squeezing which can be controlled by adjusting the system parameters. Surprisingly, the controllable simultaneous nonclassical dynamics in the same physical parameter regime which is induced by the optomechanical Kerr nonlinearity to the parametric DCE cannot occurs in the standard DCE or Kerr-type systems. The proposed nonclassical microwave radiation source possesses the potential to be applied in quantum information processing, quantum computing as well as microwave quantum sensing.
@@ I Introduction
It is well-known that quantum parametric processes provide essential mechanisms for generating and manipulating quantum states in diverse aspects of quantum science and technologies. The most remarkable property of these processes concerns the dynamical amplification of quantum vacuum fluctuations that is responsible for the creation of pairs of real particles as a consequence of strong non-adiabatic change of a system parameter or boundary condition [1, 2]. This effect, which is generally referred to as the dynamical Casimir effect (DCE) [1, 3, 4, 5, 6], describes a process in which a cavity with periodically oscillating mirrors produces pairs of photons out of the electromagnetic vacuum fluctuations. Along with the theoretical studies on the issue of the possibility of particle creation via the DCE in a large variety of systems, ranging from cosmology [7, 8, 9] to non-stationary cavity QED [10, 11, 12, 13, 14, 15], various theoretical schemes for practical applications of the DCE have been suggested, including generation of photons with nonclassical properties [16, 143, 144], generation of squeezed [19], and entangled atomic states [20], generation of multipartite entanglement in networks of quantum cavities and superconducting devices [21], generation of entanglement between two moving qubits [22], generation of EPR quantum steering and Gaussian interferometric power [23], and quantum synchronization of two qubits [24]. In addition, in the context of quantum thermodynamics, it has been shown that the DCE can be used for realizing the so-called anti-DCE [25], i.e., coherent annihilation of photons instead of creation, as a resource for work extraction from the atom-field system [26], for implementation of a quantum Otto heat engine [27], and for cooling down the cavity wall in the presence of a non-vanishing temperature gradient between the wall and the cavity [28].
@@
From an experimental point of view, for a measurable flux of real photons to be generated, the moving boundaries should oscillate at very high frequencies (more than 10GHz) that are not yet experimentally achievable. Consequently, alternative schemes based on imitation of boundary motion have been proposed, with examples including periodic modulation of the optical properties of the boundary [29, 30, 31] or of the optical path length of a cavity [32, 33, 34], amplification of the DCE in a cavity within the Rabi model in the ultrastrong coupling regime via quantum optimal control strategies [35], and time modulation of the Kerr or higher-order nonlinearities in a cavity [36]. Some other experimental schemes aiming at the observation of the DCE can be found in Refs. [29, 37, 38]. Notably, it has been reported about a successful implementation of DCE in superconducting circuits through two independent experiments by fast-modulating either the electrical boundary condition of a transmission line [39] or the effective speed of light in a Josephson metamaterial [40]. Moreover, recently an analogue DCE has been experimentally realized in the near infrared regime using a dispersion-oscillating photonic crystal fiber [41]. Results of this experiment implied that the generated DCE photons exhibit nonclassical photon anti-bunching statistics.
@@
Besides the photonic DCE, analog models for the generation of other particles than photons via the DCE have also been investigated such as the DCE of phonons in a time-modulated atomic Bose-Einstein condensate (BEC) [42, 43], emission of Bogoliubov excitations in an exciton-polariton condensate by an ultrashort laser pulse [44], phononic DCE in a time-modulated quantum fluid of light [45], DCE of magnon excitation in a spinor BEC driven by a time-dependent magnetic field [46], and DCE of phonons in a gas of laser-cooled atoms with time-dependent effective charge [47].
@@
On the other hand, during the recent decades optomechanical systems (OMSs) where electromagnetic radiation pressure is linearly or quadratically coupled to a mechanical oscillator (MO) [48] have been widely employed in a large variety of applications, for example, displacement and force sensing [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59], optomechanical cooling [60, 61, 62], quantum correlations such as entanglement [63, 64, 65, 66, 67] and synchronization [68, 69, 70], generation of nonclassical states of the mechanical and optical modes [71, 72, 73, 74, 75, 76], quadrature squeezing/amplification [77, 78, 79, 80], optomechanically induced transparency [81, 82, 83, 84, 85, 86, 87, 88, 89, 90], and quantum illumination radar [91, 92, 93].
@@
In recent years, a variety of theoretical studies have identified the OMSs as promising platforms for realizing the DCE. To mention some of them, one could highlight the DCE of phonons in a non-stationary quantum-well assisted optomechanical cavity [94]; the DCE of phononic excitation in the so-called membrane-in-the-middle optomechanical cavity [95]; the DCE of photons and mechanical-/Bogoliubov-type phonons in a modulated hybrid optomechanical cavity containing an atomic BEC [96]; the DCE of photons in an optomechanical cavity under incoherent mechanical excitation [97]; the DCE of photons in a squeezed-cavity-assisted OMS [98]; the DCE of photons in an optomechanical cavity interacting with a one-dimensional photonic crystal [99]; the DCE of phonons in the ultrastrong-coupling regime of optomechanics [100]; and the DCE of photons in non-perturbative coupling regime of cavity optomechanics [101]. Remarkably, in Ref. [101] the authors have found that the standard resonance condition for the generation of DCE photons which requires the mechanical frequency to be, at least, twice the first cavity mode frequency does not need to be satisfied, provided that the coupling between the cavity field and the moving mirror is non-vanishing with respect to the optical and mechanical resonance frequencies. In addition, they have shown that the non-perturbative regime of DCE can give rise to the steady-state entanglement between the moving mirror and the cavity field.
@@
One of the most important characteristics of cavity OMSs is a type of intrinsic nonlinearity resulting from the radiation pressure coupling. This nonlinearity stems from the fact that in a typical cavity OMS the position of the MO modulates the resonance frequency of the cavity mode. In other words, the optical length of the cavity depends on the intensity of the intracavity field and, consequently, the optomechanical cavity can behave effectively as a nonlinear Kerr medium [102, 103]. This inherent nonlinearity enables pondermotive squeezing of the cavity field [104, 105], generation of mechanical cat states [106, 107], intracavity cat-state generation [108] photon blockade [109, 110], optical bistability [111, 112], phonon–photon entanglement in the bistable regime [113], and cooling an optomechanical resonator into a cat state of motion [114].
@@
The above-mentioned cavity optomechanical phenomena, such as mechanical ground-sate cooling, entanglement, quadrature squeezing/amplification, and optomechanically induced transparency emerge in the so-called normal dissipation regime, in which the decay rate of the MO is typically significantly smaller compared to that of the cavity mode. However, recent investigations in the field of dissipation engineering have introduced a novel regime of dissipation in optomechanics, the so-called reversed-dissipation regime (RDR), where the roles of the cavity mode and the MO are interchanged [115], i.e., the dissipation rate of the MO prominently exceeds the cavity linewidth. In this regime, the MO plays the advantageous role of an additional cold dissipative reservoir for the cavity mode [116, 117]. Recently significant development  in experimental implementations has made possible the realization of the RDR in various experimental setups, including superconducting circuits [118, 119], microwave-cavity optomechanical systems [120], and mechanical resonators embedded by erbium ions [121]. In addition, a variety of applications have been proposed and investigated for this regime. Some examples include quantum-limited amplification and self-oscillation of photons [115, 120], entangled-photon generation [116], nonreciprocal optomechanically induced transparency and enhancing the ground-state cooling of a MO [122], and broadband nonreciprocal and chiral photon transmission [123].
@@
Inspired by the above-mentioned investigations, in this paper, we propose a protocol to study the parametric DCE in an experimentally realized microwave cavity OMS [120] which is dispersively driven and operates in the RDR. In such case, the degrees of freedom of the MO can be adiabatically eliminated from the system dynamics. We show that the coherent time modulation of the driving laser frequency leads to the parametric amplification of the quantum vacuum fluctuations of the intracavity field mode, resulting in the creation of microwave Casimir photons over time scales longer than the mechanical decoherence time. In addition, the adiabatic elimination of the mechanical mode will induce an effective nonlinear Kerr-type photon–photon interaction, which provides the feasibility of controllable manipulation of the quantum properties of the generated DCE photons. We analyze both analytically and numerically the effects of the induced Kerr nonlinearity as well as the modulation amplitude of the driving laser frequency on the dynamical behaviors of the mean number of the generated microwave Casimir photons, photon counting statistics, and photon quadrature squeezing. Within the short-time approximation, i.e., over time scales much shorter than the cavity decay time, we find a closed analytical expression for the system unitary time-evolution operator by using the Wei-Norman theorem [124], which is of the form of a generalized squeezing operator affected by the induced Kerr nonlinearity. The presence of the Kerr nonlinearity causes that the time evolution of the mean number of Casimir photons exhibits an oscillatory behavior whose amplitude decreases with increasing time. Regarding the nonclassical features of the generated DCE photons, we find that they can exhibit quadrature squeezing and sub-Poissonian statistics in the course of time evolution.
Beyond these general dynamical features, our work identifies several aspects that distinguish this scheme from previous DCE studies. We show that the generated Casimir photons can simultaneously exhibit sub-Poissonian statistics and Wigner-function negativity, a combination of nonclassical signatures not previously reported in DCE settings. Moreover, the proposed DCE+Kerr mechanism relies on experimentally accessible circuit-optomechanical parameters in the RDR, ensuring practical implementability. Finally, the interplay between the parametric modulation and the optomechanically-induced Kerr nonlinearity gives rise to additional dynamical signatures such as oscillatory and saturation behaviour in photon generation and periodic redistribution of squeezing between quadratures demonstrating quantum features beyond those achievable in standard DCE or fixed-Kerr models.
@@
The paper is structured as follows. First, we introduce the physical system under consideration in Sec. II, and derive an effective system Hamiltonian in the RDR of cavity optomechanics when the cavity is dispersively driven by an external modulated field. The effects of the induced Kerr nonlinearity on the dynamics of the mean number, the quantum statistics, and the quadrature squeezing of the generated Casimir photons are, respectively, discussed in Secs. III, IV and V. In Sec. VI we explore the signature of the nonclassical characteristic of the generated Casimir photons by calculating the corresponding Wigner function. In particular, we examine the impact of system dissipation on the temporal behavior of the Wigner negativity. A detailed discussion of the experimental requirements for implementation of the model is undertaken in Sec. VII. We summarize the conclusions of our work in Sec. VIII. Some detailed mathematical calculations can be found in the appendices.
@@ II Physical Model and The Effective System Hamiltonian
Figure 1: (Color online) (a) Schematic diagram of the considered microwave optomechanical circuit [120] composed of two inductively coupled microwave LC resonators which are inductively coupled to a microwave feedline as a common bath. The two hybridized dark and bright modes of the circuit that are, respectively, used as the primary $\hat{a}$ and auxiliary $\hat{a}_{\rm aux}$ interact with the vibration of the suspended top electrode of a shared capacitor, acting as a MO with resonance frequency $\omega_{m}$. The auxiliary (primary) mode is driven by a classical laser field of frequency $\omega_{L}^{\rm aux}(\omega_{L})$ and amplitude $E_{L}^{\rm aux}(E_{L})$. (b) To prepare a cold, dissipative mechanical reservoir for microwave photons the dissipation rate $\gamma_{m}$ of the MO should be increased to match the much larger dissipation rate $\kappa$ of the primary mode having resonance frequency $\omega_{c}$. For this purpose, the auxiliary mode with resonance frequency $\omega_{\rm aux}$ and dissipation rate $\kappa_{\rm aux}$ is used to damp out the MO via optomechanical sideband cooling and, hence, prepare it as a strongly dissipative, cold reservoir for the primary mode. (c) A schematic overview of the process to derive Hamiltonian (26).
As schematically illustrated in Fig. 1(a), the system we have chosen to consider is a microwave-cavity optomechanical circuit, which has been previously presented in Ref. [120] aiming at realizing a cold, dissipative mechanical reservoir for microwave photons. It is composed of two inductively coupled microwave LC resonators, both coupled inductively to a common microwave feedline, serving as a common bath, with external coupling strengths $\kappa_{1}^{\rm ex}$ and $\kappa_{2}^{\rm ex}$. One of the LC resonators contains a mechanically compliant capacitor that can be modeled as a single-mode MO. By making use of an electromagnetic mode, referred to as the auxiliary mode, the MO is damped out via optomechanical sideband cooling [125] in order to prepare it as a strongly dissipative, cold reservoir for another electromagnetic mode which referred to as the primary mode.
In order to accomplish the RDR in this setup, the auxiliary mode needs to be cooled down with a rate much faster than the electromagnetic decay rate of the primary mode, requiring the microwave cavities to have extremely different decay rates. One successful approach to overcome this challenge is to engineer hybridized modes with intrinsically different decay rates arising from interference in the output channel [for details, see the Supplementary Information in Ref. [120]]. The resulting interaction Hamiltonian is given by [120] ($\hbar=1$)
@@
where $\hat{b}$ is the annihilation operator for the mechanical mode with frequency $\omega_{m}$ and decay rate $\gamma_{m}$, $\hat{a}_{1}$ and $\hat{a}_{2}$ are the annihilation operators for the bare modes with frequency $\omega_{0}$, $J$ denotes the intermode coupling strength, and $g_{0}^{\prime}$ stands for the vacuum electromechanical coupling strength to the first bare mode. By introducing the symmetric and antisymmetric superpositions of the bare modes $\hat{a}_{s,a}=(\hat{a}_{1}\pm\hat{a}_{2})/\sqrt{2}$ [120] the intermode coupling term can be diagonalized, and thus the interaction Hamiltonian is expressed as
@@
The above Hamiltonian is written in the limit of large intermode coupling compared to the mechanical frequency, $J\gg\omega_{m}$, which allows to neglect the cross terms $\hat{a}_{s}^{\dagger}\hat{a}_{a}(\hat{b}+\hat{b}^{\dagger})$ as well as the other non-resonance terms. Although the bare modes are degenerate, the primary and auxiliary eigenmodes with respective energies of $\omega_{c}=\omega_{0}-J$ and $\omega_{\rm aux}=\omega_{0}+J$ have an energy difference of $2J$ for $J\neq 0$. As is evident from the Hamiltonian (2), the intermode coupling $(J\neq 0)$ results in that the eigenmodes are both coupled to the MO with coupling strength $g_{0}^{\prime}/2$.
The hybridized modes $\hat{a}_{s,a}$ constitute a strongly coupled (bright) and a weakly coupled (dark) mode with the respective decay rates $\kappa_{s}^{\rm ex}=\kappa_{1}^{\rm ex}+\kappa_{2}^{\rm ex}$ and $|\kappa_{a}^{\rm ex}=\absolutevalue{\kappa_{1}^{\rm ex}-\kappa_{2}^{\rm ex}}$, which can be explained as the result of the interference of the bare-mode external coupling rates $\kappa_{1}^{\rm ex}$ and $\kappa_{2}^{\rm ex}$ with the output channel [see Ref. [120] and its Supplementary Information for details]. In the case of nearly the same coupling rates for the bare modes ($\kappa_{1}^{\rm ex}\approx\kappa_{2}^{\rm ex}$) we have $\kappa_{s}^{\rm ex}\gg\kappa_{a}^{\rm ex}$ so that the antisymmetric (dark) mode is approximately decoupled from the common reservoir. Following Ref. [120], in subsequent discussions we refer the dark and the bright modes as the primary and the auxiliary modes, respectively, with respective resonance frequencies $\omega_{c}$ and $\omega_{\rm aux}$ and decay rates $\kappa$ and $\kappa_{\rm aux}$ [see Fig. 1(b)].